The SOS Reference Manual

Edition 2.4 March 1998

by Chris Hanson

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993-98 Massachusetts Institute of Technology

This material was developed by the Scheme project at the Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science. Permission to copy this document, to redistribute it, and to use it for any purpose is granted, subject to the following restrictions and understandings.

1. Any copy made of this document must include this copyright notice in full.

2. Users of this document agree to make their best efforts (a) to return to the MIT Scheme project any improvements or extensions that they make, so that these may be included in future releases; and (b) to inform MIT of noteworthy uses of this document.

3. All materials developed as a consequence of the use of this document shall duly acknowledge such use, in accordance with the usual standards of acknowledging credit in academic research.

4. MIT has made no warrantee or representation that the contents of this document will be errorfree, and MIT is under no obligation to provide any services, by way of maintenance, update, or otherwise.

5. In conjunction with products arising from the use of this material, there shall be no use of the name of the Massachusetts Institute of Technology nor of any adaptation thereof in any advertising, promotional, or sales literature without prior written consent from MIT in each case.

1 An Object System

sos is a Scheme object system derived from Tiny clos^{1}, which in turn was loosely derived from clos, the Common Lisp Object System. Its basic design and philosophy is closely related to Tiny clos, but there are differences in naming and interface.

This document is a reference manual, and as such does not attempt to teach the reader about object-oriented programming. It is assumed that you already have a passing familiarity with clos and with Scheme.

In the procedure descriptions that follow, certain argument names imply restrictions on the corresponding argument. Here is a table of those names. The parenthesised name in each entry is the name of the predicate procedure that the argument must satisfy.

class The argument must be a class (class?).

instance The argument must be an instance (instance?).

name The argument must be a symbol (symbol?); sometimes this is also allowed to be #f

          (false?).

generic-procedure The argument must be a generic procedure (generic-procedure?).

method The argument must be a method (method?).

specializer The argument must be a method specializer (specializer?).

procedure The argument must be a procedure (procedure?).

slot The argument must be a slot descriptor (slot-descriptor?).

1.1 Classes

A class is an object that determines the structure and behavior of a set of other objects, which are called its instances. However, in this document, the word instance usually means an instance of the class <instance>.

1 Tiny clos was written by Gregor Kiczales of Xerox parc; sos is derived from version 1.2 of Tiny clos.

A class can inherit structure and behavior from other classes. A class whose definition refers to other classes for the purpose of inheriting from them is said to be a subclass of each of those classes. The classes that are designated for purposes of inheritance are said to be superclasses of the inheriting class.

A class can have a name. The procedure class-name takes a class object and returns its name. The name of an anonymous class is #f.

A class C[1] is a direct superclass of a class C[2] if C[2] explicitly designates C[1] as a superclass in its definition. In this case, C[2] is a direct subclass of C[1]. A class C[n] is a superclass of a class C[1] if there exists a series of classes C[2], …, C[n-1] such that C[i+1] is a direct superclass of C[i] for all i between 1 and n. In this case, C[1] is a subclass of C[n]. A class is considered neither a superclass nor a subclass of itself. That is, if C[1] is a superclass of C[2], then C[1] is different from C[2]. The set of classes consisting of some given class C along with all of its superclasses is called “C and its superclasses.”

Each class has a class precedence list, which is a total ordering on the set of the given class and its superclasses. The total ordering is expressed as a list ordered from the most specific to the least specific. The class precedence list is used in several ways. In general, more specific classes can shadow, or override, features that would otherwise be inherited from less specific classes. The method selection and combination process uses the class precedence list to order methods from most specific to least specific.

When a class is defined, the order in which its direct superclasses are mentioned in the defining form is important. Each class has a local precedence order, which is a list consisting of the class followed by its direct superclasses in the order mentioned in the defining form.

A class precedence list is always consistent with the local precedence order of each class in the list. The classes in each local precedence order appear within the class precedence list in the same order. If the local precedence orders are inconsistent with each other, no class precedence list can be constructed, and an error is signalled.

Classes are organized into a directed acyclic graph. There are two distinguished classes, named <object> and <instance>. The class named <object> has no superclasses. It is a superclass of every class except itself. The class named <instance> is a direct subclass of <object> and is the base class for instance objects. Instances are special because sos has efficient mechanisms for dispatching on them and for accessing their slots.

1.1.1 Class Datatype

The procedures in this section may be used to construct and inspect classes.

Procedure make-class name direct-superclasses direct-slots Creates and returns a new class object.

Name is used for debugging: it is a symbol that appears in the printed representation of the class and has no role in the semantics of the class. Alternatively, name may be #f to indicate that the class is anonymous.

Direct-superclasses must be a list of class objects. The new class inherits both methods and slots from the classes in this list. Specifying the empty list for direct-superclasses is equivalent to specifying (list <instance>).

Direct-slots describes additional slots that instances of this class will have. It is a list, each element of which must have one of the following forms:

name

(name . plist)

where name is a symbol, and plist is a property list. The first of these two forms is equivalent to the second with an empty plist.

Each of the elements of direct-slots defines one slot named name. Plist is used to describe additional properties of that slot. The following properties are recognized:

initial-value

This property specifies the default initial value for the slot, i.e. the value stored in the slot when an instance is created and no value is explicitly specified by the instance constructor. If neither the initial-value nor the initializer property is specified, the slot has no default initial value.

initializer

This property specifies a procedure of no arguments that is called by an instance constructor whenever an instance containing this slot is created. The value returned by the initializer procedure is the initial value of the slot. accessor This property specifies a generic procedure; make-class will add an ac-

cessor method for this slot to the procedure. See Section 1.3 [Slots], page 14.

modifier This property specifies a generic procedure; make-class will add a modifier method for this slot to the procedure. See Section 1.3 [Slots], page 14.

initpred This property specifies a generic procedure; make-class will add an “initialized?” predicate method for this slot to the procedure. See Section 1.3 [Slots], page 14.

Slot properties are combined in slightly complicated ways.

• It is not allowed to specify both initial-value and initializer for a slot in a given call to make-class; at most one of these properties may be given.

• If a slot is specified for a given class, and a slot of the same name is inherited from a superclass, then the slot properties for the two slots are combined. Slot properties from the subclass shadow those of the superclass. However, if a superclass has a slot property, and the subclass does not, the property is inherited. The resulting class never has more than one slot of a given name.

• When combining superclass and subclass slots, initial-value and initializer shadow one another. In other words, regardless of the inherited slot properties, the resulting slot has at most one of these two properties.

Examples of make-class:

 (define <cell>
   (make-class '<cell> '() '()))
 (define <contact>
   (make-class '<contact>
               (list <cell>)
               `((name accessor ,cell-name))))
 (define <compound-cell>
   (make-class '<compound-cell>
               (list <cell>)
               `((width accessor ,cell-width)
                 (height accessor ,cell-height)
                 (components accessor ,cell-components
                             modifier ,set-cell-components!
                             initial-value ()))))

Syntax define-class name direct-superclasses direct-slot … Define name to be a class. In its basic form, define-class might have been defined by

     (define-syntax define-class
       (syntax-rules ()
         ((define-class name (class ...) slot ...)
          (define name
            (make-class (quote name)
                        (list class ...)
                        (quote (slot ...)))))))

Note that slot properties are handled specially by define-class. If a direct-slot specifies a slot properties property list, the keys of the property list (i.e. the even-numbered elements) are not evaluated, while the datums of the property list are evaluated. The expansion above does not show the proper treatment of slot properties.

In addition to the slot properties recognized by make-class, define-class recognizes a special slot property, called define. The define property specifies that some or all of the slot accessors should be defined here; that is, generic procedures should be constructed and bound to variables, and then the accessor methods added to them.

The argument to the define property is a list containing any combination of the symbols accessor, modifier, and initpred. As an abbreviation, the argument may be one of these symbols by itself, which is equivalent to the list containing that symbol. Also, the argument may be the symbol standard, which is equivalent to (accessor

modifier).

The argument to define specifies the accessors that will be defined by this form. The accessors are defined using default names, unless the names are overridden by the corresponding slot property. The default names for a class <foo> and a slot bar are

foo-bar, set-foo-bar!, and foo-bar-initialized?, respectively for the accessor,

modifier, and initpred. For example,

     (define-class foo
       (bar define accessor))

defines an accessor called foo-bar, but

     (define-class foo
       (bar define accessor accessor foo/bar))

instead defines an accessor called foo/bar. Finally,

(define-class foo
(bar accessor foo/bar))

doesn’t define any accessor, but assumes that foo/bar is a previously-defined generic procedure and adds an accessor method to it.

define-class permits the specification of class options, which are options that pertain to the class as a whole. Class options are specified by overloading name: instead of a symbol, specify a pair whose car is a symbol and whose cdr is an alist. The following class options are recognized:

(predicate [name])

Specifies that a predicate procedure should be defined for this class. Name must be either a symbol or #f: a symbol specifies the name that will be bound to the predicate procedure, and #f specifies that no predicate procedure should be defined. If name is omitted, or if no predicate option is specified, a predicate procedure is defined by appending ? to the name of the class. If the class name is surrounded by angle brackets, they are stripped off first. For example, the default predicate name for the class

<foo> is foo?.

(constructor [name] slot-names [n-init-args])

Specifies that a constructor procedure should be defined for this class. Name must be a symbol, which is the name that will be bound to the constructor procedure; if omitted, a default name is formed by prepending make- to the name of the class. If the class name is surrounded by angle brackets, they are stripped off first. For example, the default constructor name for the class <foo> is make-foo. Slot-names and n-init-args correspond to the arguments of the respective names accepted by instance-constructor, and can take any of the allowed forms for those arguments.

(separator string)

Specifies how names for slot accessors are constructed. If this option isn’t given, the name of a slot accessor is formed by concatenating the name of the class with the name of the slot, with a hyphen between them. When this option is given, string is used instead of the hyphen. For example, normally a slot accessor for the slot bar in the class foo is called foo-bar. A class option (separator ".") will cause the slot accessor to be called foo.bar, the modifier to be called set-foo.bar!, and the initialization predicate to be called foo.bar?.

Examples of define-class (compare these to the similar examples for make-class):

     (define-class <cell> ())
     (define-class (<contact> (constructor (name) no-init)) (<cell>)
       (name accessor cell-name))
     (define-class (<compound-cell> (constructor ())) (<cell>)
       (width accessor cell-width)
       (height accessor cell-height)
       (components accessor cell-components
                   modifier set-cell-components!
                   initial-value '()))

Procedure make-trivial-subclass superclass1 superclass2 … This convenience procedure makes a subclass that defines no new slots, and that inherits from the given superclasses. It is equivalent to the following

     (make-class (class-name superclass1)

(list superclass1 superclass2 …)

                 '())

Procedure class? object

Returns #t if object is a class, otherwise returns #f.

Procedure subclass? class specializer Returns #t if class is a subclass of specializer, otherwise returns #f. If specializer is a class, the result follows from the above definition of subclass, except that a class is a subclass of itself. If specializer is a record type, it is equivalent to having used the record-type-class of the record type. Finally, if specializer is a union specializer, subclass? is true if class is a subclass of one or more of the component classes of specializer.

Procedure object-class object

Returns the class of object. Object may be any Scheme object; if object is known to

be an instance, instance-class is faster than object-class.

Procedure class-name class

Returns the name of class. This is the name argument passed to make-class when class was created.

Procedure class-direct-superclasses class

Returns a list of the direct superclasses of class. If a non-empty direct-superclasses argument was passed to make-class when class was created, this list is equal? to that argument. The returned value must not be modified.

Procedure class-direct-slot-names class

Returns a list of symbols that are the names of the direct slots of class. This list contains only those slots that were defined in the call to make-class that created class; it does not contain slots that were inherited. The returned value must not be modified.

Procedure class-precedence-list class

Returns a list of the superclasses of class. The order of this list is significant: it is the method resolution order. This list will always have class as its first element, and <object> as its last element. The returned value must not be modified.

1.1.2 Predefined Classes

sos provides a rich set of predefined classes that can be used to specialize methods to any of Scheme’s built-in datatypes.

Class <object>

This is the class of all Scheme objects. It has no direct superclasses, and all other classes are subclasses of this class.

Class <instance>

This is the class of instances. It is a direct subclass of <object>. The members of this class are the objects that satisfy the predicate instance?.

Class <boolean> Class <char> Class <entity> Class <pair> Class <procedure> Class <record> Class <string> Class <symbol> Class <vector>

These are the classes of their respective Scheme objects. They are all direct subclasses of <object>. The members of each class are the objects that satisfy the corresponding predicate; for example, the members of <procedure> are the objects that satisfy

procedure?.

Class <generic-procedure>

This is the class of generic procedure instances. It is a direct subclass of <procedure>.

Class <method>

This is the class of method objects. It is a direct subclass of <instance>.

Class <chained-method> Class <computed-method> Class <computed-emp>

These classes specify additional method objects with special properties. Each class is a subclass of <method>.

The following are the classes of Scheme numbers. Note that object-class will never return one of these classes; instead it returns an implementation-specific class that is associated with a particular numeric representation. The implementation-specific class is a subclass of one or more of these implementation-independent classes, so you should use these classes for specialization.

Class <number> Class <complex> Class <real> Class <rational> Class <integer>

These are the classes of the Scheme numeric tower. <number> is a direct subclass of <math-object>, <complex> is a direct subclass of <number>, <real> is a direct subclass

of <complex>, etc.

Class <exact> Class <exact-complex> Class <exact-real> Class <exact-rational> Class <exact-integer>

These are the classes of exact numbers. <exact> is a direct subclass of <number>, <exact-complex> is a direct subclass of <exact> and <complex>, and in general, each is a direct subclass of preceding class and of the class without the exact- prefix.

Class <inexact> Class <inexact-complex> Class <inexact-real> Class <inexact-rational> Class <inexact-integer>

These are the classes of inexact numbers. <inexact> is a direct subclass of <number>, <inexact-complex> is a direct subclass of <inexact> and <complex>, and in general, each is a direct subclass of preceding class and of the class without the inexact- prefix.

1.1.3 Record Classes

sos allows generic procedures to discriminate on record types. This means that a record structure defined by means of make-record-type or define-structure can be passed as an argument to a generic procedure, and the generic procedure can use the record’s type to determine which method to be invoked.^{2}

2 If the type option of define-structure is used, the resulting data structure is not a record and thus cannot be used in this manner.

In order to support this, sos accepts record type descriptors in all contexts that accept classes. Additionally, every record type descriptor has an associated sos class; either the class or the record type can be used with equivalent results.

Procedure record-type-class record-type

Record-type must be a record type descriptor (in other words, it must satisfy the predicate record-type?). Returns the class associated with record-type.

Procedure record-class record

Record must be a record (in other words, it must satisfy the predicate record?). Returns the class associated with record. This is equivalent to

     (record-type-class (record-type-descriptor record))

1.1.4 Specializers

A specializer is a generalization of a class. A specializer is any one of the following:

• A class.

• A record type, which is equivalent to its associated class.

• A union specializer, which is a set of classes.

A specializer may be used in many contexts where a class is required, specifically, as a method specializer (hence the name), as the second argument to subclass?, and elsewhere.

Procedure specializer? object

Returns #t if object is a specializer, otherwise returns #f.

Procedure specializer-classes specializer

Returns a list of the classes in specializer. If specializer is a class, the result is a list of that class. If specializer is a record type, the result is a list of the record type’s class. If specializer is a union specializer, the result is a list of the component classes of the specializer.

Procedure specializer=? specializer1 specializer2 Returns #t if specializer1 and specializer2 are equivalent, otherwise returns #f. Two specializers are equivalent if the lists returned by specializer-classes contain the same elements.

Procedure union-specializer specializer …

Returns a union specializer consisting of the union of the classes of the arguments. This is equivalent to converting all of the specializer arguments to sets of classes, then taking the union of those sets.

Procedure union-specializer? object

Returns #t if object is a union specializer, otherwise returns #f.

Procedure specializers? object

Returns #t if object is a list of specializers, otherwise returns #f.

Procedure specializers=? specializers1 specializers2 Specializers1 and specializers2 must be lists of specializers. Returns #t if specializers1 and specializers2 are equivalent, otherwise returns #f. Two specializers lists are equivalent if each of their corresponding elements is equivalent.

1.2 Instances

An instance is a compound data structure much like a record, except that it is defined by a class rather than a record type descriptor. Instances are more powerful than records, because their representation is designed to support inheritance, while the representation of records is not.

Procedure instance-constructor class slot-names [n-init-args] Creates and returns a procedure that, when called, will create and return a newly allocated instance of class.

Class must be a subclass of <instance>. Slot-names must be a list of symbols, each of which must be the name of a slot in class. N-init-args will be described below.

In its basic operation, instance-constructor works much like record-constructor:

the slot-names argument specifies how many arguments the returned constructor accepts, and each of those arguments is stored in the corresponding slot of the returned instance. Any slots that are not specified in slot-names are given their initial values, as specified by the initial-value or initializer slot properties; otherwise they are left uninitialized.

After the new instance is created and its slots filled in, but before it is returned, it is passed to the generic procedure initialize-instance. Normally, initializeinstance does nothing, but because it is always called, the programmer can add methods to it to specify an initialization that is to be performed on every instance of the class.

By default, initialize-instance is called with one argument, the newly created instance. However, the optional argument n-init-args can be used to specify additional arguments that will be passed to initialize-instance.

The way this works is that the returned constructor procedure accepts additional arguments after the specified number of slot values, and passes these extra arguments to initialize-instance. When n-init-args is not supplied or is #t, any number of extra arguments are accepted and passed along. When n-init-args is an exact non-negative integer, exactly that number of extra arguments must be supplied when the constructor is called. Finally, if n-init-args is the symbol no-initialize-instance, then the constructor accepts no extra arguments and does not call initialize-instance at all; this is desirable when initialize-instance is not needed, because it makes the constructor significantly faster.

For notational convenience, n-init-args may take two other forms. First, it may be a list of symbols, which is equivalent to the integer that is the length of the list. Second, it may be the symbol no-init, which is an abbreviation for no-initialize-instance.

Note that the default method on initialize-instance accepts no extra arguments and does nothing.

Examples of instance-constructor:
     (define-class <simple-reference> (<reference>)
       (from accessor reference-from)
       (to accessor reference-to)
       (cx accessor reference-cx)
       (cy accessor reference-cy))

     (define make-simple-reference
       (instance-constructor <simple-reference>
                             '(from to cx cy)
                             'no-init))
     (define-class <simple-wirenet> (<wirenet>)
       (cell accessor wirenet-cell)
       (wires accessor wirenet-wires
              modifier set-wirenet-wires!
              initial-value '()))
     (define make-simple-wirenet
       (instance-constructor <simple-wirenet> '(cell)))

Procedure instance? object

Returns #t if object is an instance, otherwise returns #f.

Procedure instance-class instance

Returns the class of instance. This is faster than object-class, but it works only for instances, and not for other objects.

Procedure instance-of? object specializer Returns #t if object is a general instance of specializer, otherwise returns #f. This is equivalent to

     (subclass? (object-class object) specializer)

Procedure instance-predicate specializer

Returns a predicate procedure for specializer. The returned procedure accepts one argument and returns #t if the argument is an instance of specializer and #f otherwise.

1.3 Slots

An instance has zero or more named slots; the name of a slot is a symbol. The slots of an instance are determined by its class.

Each slot can hold one value. When a slot does not have a value, the slot is said to be uninitialized. The default initial value for a slot is defined by the initial-value and initializer slot properties.

A slot is said to be accessible in an instance of a class if the slot is defined by the class of the instance or is inherited from a superclass of that class. At most one slot of a given name can be accessible in an instance. Slots are accessed by means of slot-access methods (usually generated by

make-class).

1.3.1 Slot Descriptors

Slots are represented by slot descriptors, which are data structures providing information about the slots, such as their name. Slot descriptors are stored inside of classes, and may be retrieved from there and subsequently inspected.

Procedure class-slots class

Returns a list of the slot descriptors for class. This contains all slots for class, both direct slots and inherited slots. The returned value must not be modified.

Procedure class-slot class name error?

Returns the slot descriptor for the slot named name in class. If there is no such slot: if error? is #f, returns #f, otherwise signals an error of type condition-type:no-such-

slot.

Procedure slot-descriptor? object

Returns #t if object is a slot descriptor, otherwise returns #f.

Procedure slot-name slot

Returns the name of slot.

Procedure slot-class slot

Returns the class of slot. This is the class with which slot is associated. This is not necessarily the class that defines slot; it could also be a subclass of that class. If the slot was returned from class-slots or class-slot, then this class is the argument passed to that procedure.

Procedure slot-properties slot

Returns an alist of the properties of slot. This list must not be modified.

Procedure slot-property slot name default If slot has a property named name, it is returned; otherwise default is returned.

Procedure slot-initial-value? slot

Returns #t if slot has an initial value, and #f otherwise. The initial value is specified by the initial-value slot property when a class is made.

Procedure slot-initial-value slot

Returns the initial value for slot, if it has one; otherwise it returns an unspecified value. The initial value is specified by the initial-value slot property when a class is made.

Procedure slot-initializer slot

Returns the initializer for slot; the initializer is specified by the initializer slot property when a class is made. This is a procedure of no arguments that is called to produce an initial value for slot. The result may also be #f meaning that the slot has no initializer.

1.3.2 Slot Access Methods

The procedure make-class provides slot properties that generate methods to read and write slots. If an accessor is requested, a method is automatically generated for reading the value of the slot. If a modifier is requested, a method is automatically generated for storing a value into the slot. When an accessor or modifier is specified for a slot, the generic procedure to which the generated method belongs is directly specified. The procedure specified for the accessor takes one argument, the instance. The procedure specified for the modifier takes two arguments, the instance and the new value, in that order.

All of the procedures described here signal an error of type condition-type:no-such-slot if the given class or object does not have a slot of the given name.

Slot-access methods can be generated by the procedures slot-accessor-method, slotmodifier-method, and slot-initpred-method. These methods may be added to a generic procedure by passing them as arguments to add-method. The methods generated by these procedures are equivalent to those generated by the slot properties in make-class.

Procedure slot-accessor-method class name

Returns an accessor method for the slot name in class. The returned method has one required argument, an instance of class, and the specializer for that argument is class. When invoked, the method returns the contents of the slot specified by name in the instance; if the slot is uninitialized, an error of type condition-type:uninitialized-slot is signalled.

     (define-generic get-bar (object))
     (add-method get-bar
       (slot-accessor-method <foo> 'bar))

Procedure slot-modifier-method class name

Returns a modifier method for the slot name in class. The returned method has two required arguments, an instance of class and an object. The specializer for the first argument is class and the second argument is not specialized. When invoked, the method stores the second argument in the slot specified by name in the instance.

     (define-generic set-bar! (object bar))
     (add-method set-bar!
       (slot-modifier-method <foo> 'bar))

Procedure slot-initpred-method class name

Returns an “initialized?” predicate method for the slot name in class. The returned method has one required argument, an instance of class, and the specializer for that argument is class. When invoked, the method returns #t if the slot specified by name is initialized in the instance; otherwise it returns #f.

     (define-generic has-bar? (object))
     (add-method has-bar?
       (slot-initpred-method <foo> 'bar))

1.3.3 Slot Access Constructors

For convenience, and for consistency with the record-accessor procedures record-accessor and record-modifier, each of the above method-generating procedures has a corresponding accessorgenerator. Each of these procedures creates a generic procedure, adds an appropriate method to it by calling the corresponding method-generating procedure, and returns the generic procedure. Thus, for example, the following are equivalent:

(slot-accessor <foo> 'bar)
(let ((g (make-generic-procedure 1)))
  (add-method g (slot-accessor-method <foo> 'bar))
  g)

Procedure slot-accessor class name

Returns a generic procedure of one argument that is an accessor for the slot name in class. The argument to the returned procedure must be an instance of class. When the procedure is called, it returns the contents of the slot name in that instance; if the slot is uninitialized, an error of type condition-type:uninitialized-slot is signalled.

Procedure slot-modifier class name

Returns a generic procedure of two arguments that is a modifier for the slot name in class. The first argument to the returned procedure must be an instance of class, and the second argument may be any object. When the procedure is called, it modifies the slot name in the instance to contain the second argument.

Procedure slot-initpred class name

Returns a generic procedure of one argument that is an “initialized?” predicate for the slot name in class. The argument to the returned procedure must be an instance of class. When the procedure is called, it returns #t if the slot name in that instance is initialized, otherwise it returns #f.

1.3.4 Slot Access Procedures

Finally, there is another set of three procedures, which access the contents of a slot directly, given an instance and a slot name. These procedures are very slow by comparison with the above.

However, note the following. You can use these procedures in the body of a define-method special form in an efficient way. If the define-method specifies the correct number of arguments, the body of the form contains a call to one of these procedures and nothing else, and the specified slot name is quoted, the form is rewritten during macro-expansion time as a call to the corresponding method-generating procedure. For example, the following are equivalent:

     (define-method p ((v <foo>))
       (slot-value v 'bar))
     (add-method p
                 (slot-accessor-method <foo> 'bar))

Procedure slot-value instance name Returns the contents of the slot name in instance; if the slot is uninitialized, an error

     of type condition-type:uninitialized-slot is signalled.

Procedure set-slot-value! instance name object Modifies the slot name in instance to contain object.

Procedure slot-initialized? instance name

Returns #t if the slot name in instance is initialized, otherwise returns #f.

1.4 Generic Procedures

Like an ordinary Scheme procedure, a generic procedure takes arguments, performs a series of operations, and perhaps returns useful values. An ordinary procedure has a single body of code that is always executed when the procedure is called. A generic procedure has a set of multiple bodies of code, called methods, from which a subset is selected for execution. The selected bodies of code and the manner of their combination are determined by the classes of one or more of the arguments to the generic procedure.

Ordinary procedures and generic procedures are called with identical procedure-call syntax.

Generic procedures are true procedures that can be passed as arguments, returned as values, and otherwise used in all the ways an ordinary procedure may be used. In particular, generic procedures satisfy the predicate procedure?.

1.4.1 Generic Procedure Datatype

The following definitions are used to construct and inspect generic procedures.

Procedure make-generic-procedure arity [name]

Creates and returns a new generic procedure. The generic procedure requires arity arguments.

Arity may take one of the following forms. An exact positive integer specifies that the procedure will accept exactly that number of arguments. A pair of two exact positive integers specifies inclusive lower and upper bounds, respectively, on the number of arguments accepted; the cdr may be #f indicating no upper bound.

Name is used for debugging: it is a symbol that has no role in the semantics of the generic procedure. Name may be #f to indicate that the generic procedure is anonymous. If name is not specified, it defaults to #f.

Examples:

     (define foo-bar (make-generic-procedure 2))
     (define foo-baz (make-generic-procedure '(1 . 2) 'foo-baz))
     (define foo-mum (make-generic-procedure '(1 . #f)))

Syntax define-generic name lambda-list Defines name to be a generic procedure. Lambda-list is an ordinary parameter list, which is exactly like the parameter list in a lambda special form. This expands into

     (define name
       (make-generic-procedure arity
                               (quote name)))

where arity is determined from lambda-list.

Examples (compare to examples of make-generic-procedure):

     (define-generic foo-bar (x y))
     (define-generic foo-baz (x #!optional y))
     (define-generic foo-mum (x . y))

Procedure generic-procedure? object

Returns #t if object is a generic procedure, otherwise returns #f. Note that every generic procedure satisfies the predicate procedure?.

Procedure generic-procedure-arity generic-procedure

Returns the arity of generic-procedure, as specified in the call to make-generic-procedure. The returned arity must not be modified.

Procedure generic-procedure-name generic-procedure

Returns the name of generic-procedure, as specified in the call to make-generic-

procedure.

1.4.2 Method Storage

Methods are stored in generic procedures. When a generic procedure is called, it selects a subset of its stored methods (using method-applicable?), and arranges to invoke one or more of the methods as necessary. The following definitions provide the means for adding methods to and removing them from a generic procedure.

Procedure add-method generic-procedure method Adds method to generic-procedure. If generic-procedure already has a method with the same specializers as method, then the old method is discarded and method is used in its place.

Procedure delete-method generic-procedure method Removes method from generic-procedure. Does nothing if generic-procedure does not contain method.

Procedure add-methods generic-procedure methods Adds methods, which must be a list of methods, to generic-procedure. Equivalent to calling add-method on each method in methods.

Procedure generic-procedure-methods generic-procedure

Returns a list of the methods contained in generic-procedure. The returned list must not be modified.

1.4.3 Effective Method Procedure

When a generic procedure is called, it arranges to invoke a subset of its methods. This is done by combining the selected methods into an effective method procedure, or emp, then tail-recursively invoking the emp. compute-effective-method-procedure is the procedure that is called to select the applicable methods and combine them into an emp.

Procedure compute-effective-method-procedure generic-procedure classes

Collects the applicable methods of generic-procedure by calling method-applicable? on each method and on classes. Combines the resulting methods together into an effective method procedure, and returns that emp.

Procedure compute-method generic-procedure classes This procedure is like compute-effective-method-procedure, except that it returns the result as a method whose specializers are classes.

     compute-method is equivalent to
          (make-method classes
                       (compute-effective-method-procedure generic-procedure

classes))

1.5 Methods

A method contains a method procedure and a sequence of parameter specializers that specify when the given method is applicable.

A method is not a procedure and cannot be invoked as a procedure. Methods are invoked by the effective method procedure when a generic procedure is called.

1.5.1 Method Datatype

The following procedures are used to construct and inspect methods.

Procedure make-method specializers procedure Creates and returns a new method. Note that specializers may have fewer elements than the number of required parameters in procedure; the trailing parameters are considered to be specialized by <object>.

After the returned method is stored in a generic procedure, Procedure is called by the effective method procedure of the generic procedure when the generic procedure is called with arguments satisfying specializers. In simple cases, when no method combination occurs, procedure is the effective method procedure.

Procedure method? object

Returns #t iff object is a method, otherwise returns #f.

Generic Procedure method-specializers method

Returns the specializers of method. This list must not be modified.

Generic Procedure method-procedure method

Returns the procedure of method.

Procedure method-applicable? method classes

This predicate is used to determine the applicability of method. When a method is contained in a generic procedure, and the procedure is applied to some arguments, the method is applicable if each argument is an instance of the corresponding method specializer, or equivalently, if each argument’s class is a subclass of the corresponding method specializer.

method-applicable? determines whether method would be applicable if the given arguments had the classes specified by classes. It returns #t if each element of classes is a subclass of the corresponding specializer of method, and #f otherwise.

1.5.2 Method Syntax

The following syntactic forms greatly simplify the definition of methods, and of adding them to generic procedures.

Syntax define-method generic-procedure lambda-list body … Defines a method of generic-procedure. Lambda-list is like the parameter list of a lambda special form, except that the required parameters may have associated specializers. A parameter with an associated specializer is written as a list of two elements: the first element is the parameter’s name, and the second element is an expression that evaluates to a class.

Lambda-list must contain at least one required parameter, and at least one required parameter must be specialized.

A define-method special form expands into the following:

(add-method generic-procedure

       (make-method (list specializer . . .)
         (lambda (call-next-method . stripped-lambda-list)

body …)))

where stripped-lambda-list is lambda-list with the specialized parameters replaced by their names, and the specializers are the corresponding expressions from the specialized parameters. If necessary, the specializers are interspersed with references to <object> in order to make them occur in the correct position in the sequence.

For example,

     (define-method add ((x <integer>) (y <rational>)) ...)

expands into

     (add-method add
       (make-method (list <integer> <rational>)
         (lambda (call-next-method x y) ...)))

Note that the list of specializers passed to make-method will correspond to the required parameters of the method; the specializer corresponding to a non-specialized required parameter is <object>.

Further note that, within the body of a define-method special form, the free variable call-next-method is bound to a “call-next-method” procedure (see make-chained-method for details). If the define-method body refers to this variable, the defined method is a chained method, otherwise it is an ordinary method.

Syntax method lambda-list body … This special form evaluates to an anonymous method in the same way that lambda evaluates to an anonymous procedure. Lambda-list and body have exactly the same syntax and semantics as the corresponding parts of define-method. Note that the following are completely equivalent:

     (define-method g ((a <foo>) b)
       (cons b a))
     (add-method g
       (method ((a <foo>) b)
         (cons b a)))
     (add-method g
       (make-method (list <foo>)
         (lambda (a b)
           (cons b a))))

1.5.3 Chained Methods

Sometimes it is useful to have a method that adds functionality to existing methods. Chained methods provide a mechanism to accomplish this. A chained method, when invoked, can call the method that would have been called had this method not been defined: it is passed a procedure that will call the inherited method. The chained method can run arbitrary code both before and after calling the inherited method.

Procedure make-chained-method specializers procedure Create and return a chained method. Procedure must be a procedure of one argument that returns a procedure. When the chained method is combined, its procedure will be called with one argument, a “call-next-method” procedure; it must then return another procedure that will be called when the method is invoked. The “call-next-method” procedure may called by the method procedure at any time, which will invoke the next less-specific method. The “call-next-method” procedure must be called with the same number of arguments as the method procedure; normally these are the same arguments, but that is not required.

Procedure chained-method? object

Returns #t if object is a chained method, otherwise returns #f. Note that every chained method satisfies method?.

1.5.4 Computed Methods

A computed method is a powerful mechanism that provides the ability to generate methods “on the fly”. A computed method is like an ordinary method, except that its procedure is called during method combination, and is passed the classes of the arguments in place of the arguments themselves. Based on these classes, the computed method returns an ordinary method, which is combined in the usual way.

Note that computed methods and computed emps both satisfy the predicate method?. They are not really methods in that they cannot be combined with other methods to form an effective method procedure; however, they are treated as methods by procedures such as add-method and

method-specializers.

Procedure make-computed-method specializers procedure Create and return a computed method. Procedure will be called during method combination with the classes of the generic-procedure arguments as its arguments. It must return one of the following:

• An ordinary method (as returned by make-method or make-chained-method). The returned method’s specializers must be restrictions of specializers, i.e. each specializer in the returned method must be a subclass of the corresponding specializer in specializers. In the usual case, the returned method’s specializers are the same as specializers.

• A procedure, which is converted into an ordinary method by calling make-method on specializers and the returned procedure.

• #f, which means that the computed method declines to generate a method.

Procedure computed-method? object

Returns #t if object is a computed method, otherwise returns #f.

A computed emp takes the computed-method mechanism one step further. A computed emp is like a computed method, except that it returns an effective method procedure rather than a method. compute-effective-method-procedure tries each of the applicable computed emps, and if exactly one of them returns an emp, that is the resulting effective method procedure.

Procedure make-computed-emp key specializers procedure Create and return a computed emp. Procedure will be called during method combination with the classes of the generic-procedure arguments as its arguments. It must return either an emp or #f.

Key is an arbitrary object that is used to identify the computed emp. The key is used by add-method and delete-method to decide whether two computed emps are the same; they are the same if their keys are equal?. This is necessary because a generic procedure may have more than one computed emp with the same specializers.

Procedure computed-emp? object

Returns #t if object is a computed emp, otherwise returns #f.

Generic Procedure computed-emp-key computed-emp

Returns the key for computed-emp.

1.6 Printing

The following procedures can be used to define a custom printed representation for an instance. It is highly recommended that instances be printed by write-instance-helper, as this ensures a uniform appearance for all objects.

Generic Procedure write-instance instance port

This is called by the runtime system to generate the printed representation of instance. The methods of this procedure should write the representation to port.

Procedure write-instance-helper name instance port thunk

This writes a standardized “frame” for a printed representation method. It generates the following output on port:

#[name hash-number…]

where hash-number is the result of calling hash on instance, and … is the output generated by thunk.

28 The SOS Reference Manual Variable Index 29

Variable Index

<

<boolean> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 <chained-method> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <char> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <complex> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <computed-emp> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <computed-method> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <entity> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <exact-complex> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <exact-integer> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <exact-rational> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <exact-real> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <exact> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <generic-procedure>. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <inexact-complex> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <inexact-integer> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <inexact-rational>. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <inexact-real> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <inexact> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <instance> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 <integer> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <method> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <number> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <object> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 <pair> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <procedure> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <rational> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <real> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 <record> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <string> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <symbol> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 <vector> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

A

add-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 add-methods. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

C

chained-method? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 class-direct-slot-names . . . . . . . . . . . . . . . . . . . . . . . . . . 8

                                                    class-direct-superclasses. . . . . . . . . . . . . . . . . . . . . . . . 8

class-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 class-precedence-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 class-slot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 class-slots. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 class? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

                                                    compute-effective-method-procedure . . . . . . . . . . . . 22

compute-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 computed-emp-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 computed-emp? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 computed-method? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 constructor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

D

define-class. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 define-generic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 define-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 delete-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

G

generic-procedure-arity . . . . . . . . . . . . . . . . . . . . . . . . . 21

                                                    generic-procedure-methods. . . . . . . . . . . . . . . . . . . . . . . 21

generic-procedure-name . . . . . . . . . . . . . . . . . . . . . . . . . . 21 generic-procedure?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

I

initialize-instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 instance-class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 instance-constructor . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 instance-of? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 instance-predicate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 instance? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

M

make-chained-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 make-class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 make-computed-emp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 make-computed-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

30 The SOS Reference Manual

make-generic-procedure . . . . . . . . . . . . . . . . . . . . . . . . . . 20 make-method. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 make-trivial-subclass . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 method-applicable?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 method-procedure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 method-specializers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 method? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

O

object-class. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

P

predicate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

R

record-class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 record-type-class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

S

separator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 set-slot-value! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 slot-accessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 slot-accessor-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 slot-class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 slot-descriptor? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 slot-initial-value. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 slot-initial-value? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 slot-initialized? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 slot-initializer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 slot-initpred . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 slot-initpred-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 slot-modifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 slot-modifier-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 slot-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 slot-properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 slot-property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 slot-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 specializer-classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 specializer=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 specializer? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 specializers=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 specializers? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 subclass? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

U

union-specializer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 union-specializer?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

W

write-instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 write-instance-helper . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

Concept Index 31

Concept Index

A

accessibility of slots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 accessor, for slot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

C

chained method. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 class name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 class options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 class precedence list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 computed emps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 computed method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 constructor, class option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

D

direct subclass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 direct superclass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

E

effective method procedure. . . . . . . . . . . . . . . . . . . . . . . . . 22 emp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

G

generic procedure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

I

initialize-instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1, 12

L

local precedence order. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

M

method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 modifier, for slot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

N

name, of class. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

O

order, local precedence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

P

precedence list, class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 precedence order, local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 predefined classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 predicate, class option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 printing instances . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

R

record class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

S

separator, class option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 slot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 slot accessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 slot descriptor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 slot modifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 slot, uninitialized . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 specializer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 subclass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 subclass, direct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 superclass. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 superclass, direct. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

U

uninitialized slot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

32 The SOS Reference Manual

Table of Contents

1 An Object System. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.1.1 Class Datatype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.1.2 Predefined Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.1.3 Record Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.1.4 Specializers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

1.2 Instances . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 1.3 Slots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

1.3.1 Slot Descriptors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 1.3.2 Slot Access Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 1.3.3 Slot Access Constructors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 1.3.4 Slot Access Procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

1.4 Generic Procedures. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 1.4.1 Generic Procedure Datatype. . . . . . . . . . . . . . . . . . . . . . . . . . . 19 1.4.2 Method Storage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 1.4.3 Effective Method Procedure . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

1.5 Methods. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 1.5.1 Method Datatype. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 1.5.2 Method Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 1.5.3 Chained Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 1.5.4 Computed Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

1.6 Printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

ii The SOS Reference Manual